a. Find the orthogonal and normal canonical forms of 2y^2-2yz+2zx-2xy.
b. The operation,* defined by a*b= sin(ab), is a binary operation on N
True or false with full explanation
1
Expert's answer
2021-09-14T06:08:51-0400

Answer:-

"(a) \ 2y^2-2yz+2zx-2xy"

The matrix of the quadratic equation is;

"A=\begin{vmatrix}
 0 & -1 &1\\
 -1 & 2&0\\
1&0 & -2
\end{vmatrix}"


The characteristic roots are;

i.e

"A=\begin{vmatrix}
 0-\lambda& -1 &1\\
 -1 & 2-\lambda&0\\
1&0 & -2-\lambda
\end{vmatrix}=\begin{vmatrix}
 0& -1 &1\\
 -1 & 2-\lambda&0\\
1&0 & -2-\lambda
\end{vmatrix}"


The characteristic vector for "\lambda = 0" is given by;

i.e

"-1x_2 +x_3 = 0\\
-1x_1 +2x_2 = 0\\
x_1 -2x_3 = 0"


Solving for first 2,

"\dfrac{x_1}{2}= -x_2 = x_3"


giving the Eignen vector "X_1 = k_1(2,-1,1)"


When "\lambda = 3", the characteristic vector becomes

"-4x_2 -2x_3 = 0\\
-4x_1 -1x_2 = 0\\
-2x_1 -5x_3 = 0"


which can then become

"4x_2+2x_3 = 0\\
4x_1 +1x_2 = 0\\
2x_1 +5x_3 = 0"


solving again we get

"x_1 = \dfrac{x_2}{4}= \dfrac{x_3}{8}"


"X_2 = k_2(1,4,8)"


Similarly, "X_3= k(1,8,4)"


"X_1.X_2 = X_1.X_3 = X_2.X_3= 0"


The normalised vector is

"A=\begin{vmatrix}
 -1/3& 2/3 &1/3\\
 4/3 & 1/3&8/3\\
1/3&8/3&4/3
\end{vmatrix}"


b)

"a*b = sin(ab)"

Let a be 2 and b be 5

"sin(10°) = 0.174\\
 sin(10)\ r = -0.544"


since, sinab "\neq" N,

the operation is not binary